# JEL Classifications: C61; D15; D52; E21; G51

### v2

@@ Appendix B Proof of Section 2 Results
$c_{n}(w,z)\geq b\quad\text{for all }n\geq 0,w\geq b,\text{ and }z\in\mathsf{Z}.$ (45)
To see this, recall that based on the monotonicity of $T$, we have $c_{n}\leq c_{n-1}$ for all $n\geq 1$. Recall $\bar{w}_{c}$ and $\bar{w}$ defined in (28). To simplify notation, we denote $\bar{w}_{n}:=\bar{w}_{c_{n}}$. Then based on the monotonicity of $u$, we have $\bar{w}_{n}\geq\bar{w}$ for all $n\geq 0$. Moreover, by definition, $\bar{w}(z)>0$ for all $z\in\mathsf{Z}$. Hence, we can choose $b$ such that
$b\in\left(0,\min_{z\in\mathsf{Z}}\bar{w}(z)\right).$
Then we have $c_{n}(b,z)=b$ by Proposition 2.4. Then, by the monotonicity of $c_{n}$, we have $c_{n}(w,z)\geq c_{n}(b,z)=b>0$ for all $w\geq b,z\in\mathsf{Z}$, and $n\geq 0$. This combined with (44) implies that
@@
$\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (46)
@@
$\psi_{h}=\begin{cases}1,&\text{if }h<i,\\ \alpha b^{\gamma_{i}-\gamma_{h}},&\text{if }h\geq i,\end{cases}$ (47)
@@
$\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$ (49)
@@
Appendix E Algorithms
This section presents the quantitative algorithms used in Section 4. We first describe the endogenous grid method for solving the optimal consumption policy, followed by the algorithms for computing state-dependent impulse responses.

### v3

@@ Appendix B Proof of Section 2 Results
(44) $c_{n}^{0}(w,z)\geq b\quad\text{for all }n\geq 0,w\geq b,\text{ and }z\in\mathsf{Z}.$
To see this, recall that based on the monotonicity of $T$, we have $c_{n}^{0}\leq c_{n-1}^{0}$ for all $n\geq 1$. Recall $\bar{w}_{c}$ and $\bar{w}$ defined in (27). To simplify notation, we denote $\bar{w}_{n}:=\bar{w}_{c_{n}^{0}}$. Then based on the monotonicity of $u$, we have $\bar{w}_{n}\geq\bar{w}$ for all $n\geq 0$. Moreover, by definition, $\bar{w}(z)>0$ for all $z\in\mathsf{Z}$. Hence, we can choose $b$ such that $b\in\left(0,\min_{z\in\mathsf{Z}}\bar{w}(z)\right)$. Then we have $c_{n}^{0}(b,z)=b$ by Proposition 2.4. Then, by the monotonicity of $c_{n}^{0}$, we have $c_{n}^{0}(w,z)\geq c_{n}^{0}(b,z)=b>0$ for all $w\geq b,z\in\mathsf{Z}$, and $n\geq 0$. This combined with (43) implies that
@@
(45) $\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
(46) $\psi_{h}=\begin{cases}1,&\text{if }h<i,\\ \alpha b^{\gamma_{i}-\gamma_{h}},&\text{if }h\geq i,\end{cases}$
@@
(48) $\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
@@
Appendix E Quantitative Algorithms
In this section, we present the quantitative algorithms used in Section 4. We first describe the endogenous grid method for solving the optimal consumption policy, followed by the algorithms for computing state-dependent impulse responses.